OpenStudy (anonymous):

First, factor the trinomial as a binomial square, and then use the square root method to solve. Round the solution to the nearest tenth. x2 - 14x + 49 = 3 {5.3, 8.7} {4.8, 9.2} {3.5, 7.5} {6.2, 8.3}

OpenStudy (anonymous):

(x-7)^2 = 3 (x-7) = +/- sqrt(3) then add 7 to both sides

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

answer is 5.3,8.3